23rd June 2008 APM comfy chairs 11:30-13:05 Present: RWA, MJI, JRL, PSB, MR, DWE, STH, RGM Apologies: EGS, NAW Agenda ------ 1. Actions from last meeting 2. Comments on WFAU minutes 3. Recent meetings - Video meeting U. Herts - Plato meeting ESTEC - LSST meeting Edinburgh - CACC meeting - UKIRT board meeting 4. Data archives update - AAT, ING, WFCAM, UKIRT 5. Optical/NIR processing - HawkI - IPHAS/UVEX - MegaCam/CFHTLS photometry 6. WFCAM update - illum corrs yet more tests - photometric calibration paper - data transfers - flippin' jitters (telecon?) - processing status 08A - DR4 checking - flagging rejected MSBs 7. VISTA update - general news - SADT + P2PP workshop at ESO - pipeline progress (random jitters) - data connections to ESO; ftp tests - JPE visit and Paranal camera data - survey progress web page requirements - web page technical updates 8. AOB - network upgrades - hardware - APM air con Minutes ------- 1. Actions from last meeting STH has partially checked that the Plone astrometry technical pages are correct - ongoing <<<< MJI acquired the transmission curve for the VIRCAM Z filters. He is still trying to obtain the N118 transmission curve - ongoing <<<< MJI finished off checking and acquiring the latest SMPs RGM JRL postponed reviving the changes log web page for the WFCAM pipeline until the WFCAM data had been reduced - ongoing <<<< STH has dropped the action to investigate the new SADT since there will now be an ESO workshop on the issue. MJI has obtained a GANTT chart of the latest commissioning plans. He will still poke people to get more detail when he can. STH Details about the LAS tests on the illumination correction were obtained by a different route. STH has obtained the graphic showing the size of the photometric radial distortion correction for the calibration paper, but has not added this to the Plone technical pages - ongoing <<<< MJI did not find out at the CACC meeting what the network costing model currently is MJI has ordered three small meeting tables for the APM building DWE found out from Andy Batey that the wall display screen would be more expensive than just going out and buying a flat-screen telly. Part of this was the "need" to have it installed on the wall by the company. The plan now to to just buy a television and have it free-standing on the end of the table (and against the wall). MJI will check the connections with his recently-bought one to see if displaying from his laptop works well. He will also measure the dimensions. <<<< 2. Comments on WFAU minutes 3. Recent meetings Video meeting U. Herts: MJI to check with EGS what deep-stacking work he has promised to do for Video. <<<< MJI to obtain JPE's presentation from the Video meeting. <<<< - Plato meeting ESTEC - LSST meeting Edinburgh - CACC meeting - UKIRT board meeting 4. Data archives update - AAT, ING, WFCAM, UKIRT 5. Optical/NIR processing HawkI: STH/MJI to investigate the photometry affected by nebulosity in the HawkI data. <<<< JRL to produce some HawkI publicity for CASU. This could be a single tile and presented in a web page. It should also contain the information about how good the photometric calibration is. <<<< - IPHAS/UVEX - MegaCam/CFHTLS photometry 6. WFCAM update illum corrs yet more tests: MJI to write a stand-alone programme that will apply the illumination correction. <<<< - photometric calibration paper - data transfers flippin' jitters (telecon?): JRL to decide if flipping jitters telecon with JAC is needed. <<<< - processing status 08A - DR4 checking flagging rejected MSBs: JRL to implement the flagging of rejected MSBs. <<<< (not sure if this was an action) 7. VISTA update - general news SADT + P2PP workshop at ESO: STH/EGS to attend SADT/P2PP workshop at ESO <<<< (not sure if this was an action) - pipeline progress (random jitters) - data connections to ESO; ftp tests - JPE visit and Paranal camera data - survey progress web page requirements - web page technical updates 8. AOB - network upgrades - hardware - APM air con Continuing actions ------------------ STH check that Plone astrometry technical pages are correct MJI acquire transmission curves for VIRCAM N118 filter JRL revive changes log web page for WFCAM pipeline STH include graphic showing the size of the photometric radial distortion correction in the Plone technical pages. New actions ----------- MJI check the connections on TV to see if displaying from laptop works. Also measure dimensions. MJI check with EGS what deep-stacking work has been promised for Video MJI obtain JPE's presentation from Video meeting STH investigate photometry affected by nebulosity in HawkI data MJI JRL produce web page of HawkI publicity for CASU MJI write stand-alone programme to apply illumination correction JRL decide if flipping jitters telecon with JAC is needed JRL implement flagging of rejected MSBs (not sure if this was an action) STH attend SADT/P2PP workshop at ESO EGS (not sure if this was an action)